Can someone help me on this?
I need to add the following in the lowest terms:
2/7 + 2/7 + 6/7
Thanks, Lyn

Answer by fractalier(2101)

Since all of these are like denominators, we can add numerators directly and get
2/7 + 2/7 + 6/7 = 10/7 = 1 3/7